# Adversarial thermodynamics

### v1

@@ Abstract
In thermodynamics, an agent’s ability to extract work is fundamentally constrained by their environment. Traditional frameworks struggle to capture how strategic decision-making under uncertainty — particularly an agent’s tolerance for risk — determines the trade-off between extractable work and probability of success in finite-scale experiments. Here, we develop a framework for non-equilibrium thermodynamics based on adversarial resource theories, in which work extraction is modelled as an adversarial game for an agent extracting work. Within this perspective, we recast the Szilard engine as a game isomorphic to Kelly gambling, an information-theoretic model of optimal betting under uncertainty — but with a thermodynamic utility function. Extending the framework to finite-size regimes, we apply a risk-reward trade-off to find an interpretation of the Renyi-divergences, in terms of extractable work for a given failure probability. By incorporating risk sensitivity via utility functions, we show that the guaranteed amount of work a rational agent would accept instead of undertaking a risky protocol is given by a Rényi divergence. This provides a unified picture of thermodynamics and gambling, and highlights how generalized free energies emerge from an adversarial setup.
@@ I Introduction
In recent decades, advances in nanotechnology and biophysics have made the study of non-equilibrium and small-scale thermodynamics increasingly important. In these regimes—where the thermodynamic limit no longer applies and systems contain only finitely many interacting particles—ensemble averages fail and fluctuations dominate. Deterministic quantities such as free energy must be replaced by probabilistic, protocol-dependent notions, leading to tradeoffs between extractable work and probability of success. The breakdown of determinism has motivated two main approaches to understand the second law of thermodynamics in regimes where standard assumptions break down: stochastic thermodynamics, and the resource-theoretic approach (for a review see [1]).
@@
Stochastic thermodynamics[2, 3] embraces the inherent randomness of small, out-of-equilibrium systems, by treating work as a fluctuating quantity. It describes how entropy on average must increase, while accounting for rare transient decreases. On the other hand, the resource-theoretic approach[4, 5, 6] reformulates the second law as a constraint on possible state transformations. It shifts the focus from averages to operational possibilities, yielding generalized second laws (expressed through Renyi divergences) that quantify how far a state is from thermal equilibrium[7, 8]. While both frameworks are profoundly successful, and their views on the second law can be related[9], neither provides a complete, operational prescription for how an agent’s strategic choices—specifically, their tolerance for risk—directly determine the trade-off between the amount of work they can extract and the probability of successfully obtaining it in a single, finite-scale experiment. In this work, we bridge this gap by treating work extraction from the lens of expected utility theory and decision theory [10].
@@
There have been seminal works forging connections between stochastic thermodynamics and gambling  [11, 12, 13, 14, 15, 16]. One strand of research has implemented specific gambling strategies directly within thermodynamic protocols, as seen in the work extraction experiments of [13, 14]. Concurrently, [12] established formal links between gambling, work extraction, and information flows from an information-thermodynamic perspective. In a separate approach, [16] applied expected utility theory to evaluate thermodynamic processes as lotteries with pre-defined work payoffs. While these works successfully import concepts from one field to the other, our framework provides a deeper analogy between them and reveals a decision-theoretic aspect of thermodynamics. By formally modeling work extraction as an adversarial game implemented by a Szilard-type engine, we demonstrate that generalized free energies emerge from the necessary constraints on rational, strategic play.
@@
We develop a framework based on adversarial resource theories [17], in which an agent must choose a work-extraction protocol while contending with constraints initially imposed on the engine. This perspective casts thermodynamics as a decision-theoretic problem, where the optimal strategy is determined by the agent’s sensitivity to fluctuations. The mathematical structure of this game is formally analogous to the Kelly betting problem from information theory [18], a connection we develop in a companion paper [17]. However, a crucial distinction arises from the physical context: while financial wealth in Kelly gambling grows multiplicatively, thermodynamic work is an additive quantity. This difference dictates distinct classes of rational utility functions—Constant Relative Risk Aversion (CRRA) for gambling versus Constant Absolute Risk Aversion (CARA) for thermodynamics—which in turn shape the optimal strategies and their interpretation.
@@
The paper is structured as follows. In section, II, we formulate a Szilard-type engine as an adversarial setup between an agent (Alice) extracting the work and another agent (Bob) preparing initial constraints, and derive the average extractable work and its connection to the non-equilibrium free energy. In section III, we analyse the risk-reward trade-off through the lens of decision theory, and in particular expected utility theory, where we identify Constant Absolute Risk Aversion (CARA) as the relevant utility class for thermodynamics. This framework allows us to compute both the certainty equivalent (the guaranteed amount of work a risk-sensitive individual would accept instead of undertaking a risky extraction protocol) as Eq (9), and the expected value of the work extracted by a rational individual, Eq (8) for any level of risk aversion. We find that both are parametrized by Renyi divergences. We explain why the former result has no analogue in gambling, while the latter result is related to the gambling result derived in [17]. In Section IV, we approach the problem from an information-theoretic viewpoint on the finite-size regime, where fluctuations dominate. Here, we apply results from concurrent work [19, 17] to show that the work extraction problem reduces to a decision-theoretic problem, introducing a fundamental risk-reward trade-off in which the work extracted if a strategy is successful can be bounded in terms of Renyi divergences $D_{\alpha}$. This gives an operational interpretation to $D_{\alpha}$ for each individual $\alpha$, in terms of the minimal work extraction given some risk tolerance. This extends the result of  [20] who gave an interpretation of the Renyi divergences $D_{0}$ and $D_{\infty}$ in terms of bounds on extracting work in the case of an extremely risk averse or risk seeking agent. In both the expected utility approach and the information-theoretic approach, we give identical and explicit strategies for achieving work extraction given some risk tolerance via Eq (7). Together, all these results provide a unified operational interpretation that bridges stochastic and resource-theoretic views of the second law. Technical details and extended calculations are deferred to the appendices.
@@ II The adversarial Szilard engine
We begin by formalising a thermodynamic work extraction problem as a set-up between adversaries, mirroring the structure of Kelly betting, which we review in appendix A. Consider three players: Bob, who sets initial constraints; Alice, who optimizes work extraction (analogous to a gambler allocating bets), and Charlie, a referee who enforces randomness.
We consider an empty box of volume $V$ which Bob divides into two parts by placing a partition at some position $Q_{B}$. The referee, Charlie, then samples from a binary probability distribution $P_{X}(x)$ and places a molecule on the left or right hand side of the box according to the outcome which Alice and Bob do not know. Alice proceeds to extract work by performing isothermal compressions and expansions on the box, moving the partition to a final position $Q_{A}$ of her choice. The process is illustrated in Fig. 1.
(a) Bob places the partition at some initial position $Q^{B}$
(b) Alice chooses a position $Q^{A}$ to which she moves the partition
(c) When the molecule is on the left hand side of the box, Alice extracts work by moving it to her desired position
(d) When the molecule is on the right hand side of the box, Alice must do work to move the partition to the position of her choice
@@
Without loss of generality, assume that Bob places the partition closer to the leftmost edge of the box, so that the volume on the left hand side can be expressed as $Q_{B}\textit{l}A$, and that Alice moves her partition to a position so that the volume on the left hand side can be expressed as $Q_{A}\textit{l}A$ in terms of fractions $Q_{A}\in(0,1)$ and $Q_{B}\in(0,1)$ of the total length $l$ of the box.
The work extracted by Alice depends on the molecule’s position:
  • When $x=0$ (left-hand side): $w_{0}=k_{B}T\ln(Q^{A}/Q^{B})$
  • When $x=1$ (right-hand side): $w_{1}=k_{B}T\ln\left(\frac{1-Q^{A}}{1-Q^{B}}\right)$
@@
To maintain consistency with the notation of the literature [21], we write $Q^{B}=Q_{X}^{B}(0)$ and $1-Q^{B}=Q_{X}^{B}(1)$, and similarly for Alice. With this notation, the amount of work extracted (which is a random variable) can be written in terms of the positions to which Alice and Bob move the partition (which are not a random variable).
When the process is repeated a large number $n$ of times, the average amount of work $W$ extracted by Alice is given by $W=n\bigg(P_{X}(0)\ln{\bigg(\frac{Q_{X}^{A}(0)}{Q_{X}^{B}(0)}}\bigg)+P_{X}(1)\ln\bigg(\frac{Q_{X}^{A}(1)}{Q_{X}^{B}(1)}\bigg)\bigg)k_{B}T$ i.e.
$\boxed{W=n\big(D(P_{X}||Q^{B}_{X})-D(P_{X}||Q^{A}_{X})\big)k_{B}T}$ (1)
@@
This formula is true not only for the above binary adversarial Szilard engine (Fig. 1), but also for general multi-level engines (see Appendix B for details).
@@
The quantity (1) is maximised when Alice knows the prior $P_{X}$ and sets $Q_{X}^{A}=P_{X}$, in which case $W=D(P_{X}||Q_{X}^{B})k_{B}T$. This maximal extractable work from the system can be identified with the non-equilibrium free energy from thermodynamics up to a change in the base of the logarithm. In the particular case where Bob puts the partition in the middle, one recovers Szilard’s result $(1-H(P_{X}))\kappa_{B}T\ln{2}$ [22] in terms of entropy  [^1] It is useful to remember that for an isothermal process, the change in free energy for compression or expansion is given by $\Delta F=N\kappa_{B}T\ln(V_{f}/V_{i})$ where we have used the convention that work done on the gas is negative, and work done by the gas is positive..
@@
The expression for the average extracted work by Alice is formally related to a formula appearing in the classical Kelly betting problem from information theory [18]: both involve a difference of relative entropies, and are equivalent up to an exponentiation. We develop this connection in detail in a companion paper  [^2] However, in Kelly’s gambling problem, wealth evolves multiplicatively, making the natural utility functions those of the CRRA type, as considered in [26].
@@
It is also worth noting that Eq. (1) has also appeared in the work of [25] in the context of work extraction with an incorrect prior distribution. When Alice does not know the correct prior, our adversarial Szilard engine reduces to the formulation of [25] at the level of the ensemble average. However, here we are interested in the regime of finite-size $n$ without ensemble averages, and in the case where Alice knows the prior distribution. This will allow us to connect Eq. (1) to expected utility theory, economics, and Kelly betting, all of which assume a known distribution in the ideal case. We will thus assume that the prior distribution is known to Alice. Readers unfamiliar with Kelly betting may find it helpful to read the review in appendix A before proceeding to the next section.
III Utility maximisation: connection to expected utility formulations
@@
Recall that varying levels of risk aversion are defined in expected utility theory by the willingness to compromise between amounts of wealth and uncertainty. In this sense, the optimal strategy for different agents depends on their particular level of risk aversion. A notable example connecting the Kelly utility function to other forms of risk aversion is given in [26], where the authors analyze Kelly gambling from the perspective of Constant Relative Risk Aversion (CRRA). CRRA describes a class of utility functions in which an individual’s relative risk aversion remains constant regardless of their level of wealth. For example, if a person has more wealth, they might risk a larger absolute amount while keeping the proportion of wealth they are willing to risk constant. This property makes CRRA particularly suitable for analyzing the Kelly paradigm, as it aligns with the proportional nature of Kelly betting, since it matches financial scenarios where wealth grows multiplicatively through compounding returns.
On the other hand, in thermodynamics and Szilard engines, the system responds to changes in fixed energy amounts. Losing a fixed amount of energy matters the same regardless of total system energy- this is captured in the economics literature by the CARA utility function. This utility function has already proven useful in other thermodynamic contexts [16]. The distinction between the two types of risk aversion reflects how thermodynamic systems respond to additive shocks, whilst financial systems experience multiplicative shocks (where gains and losses compound like interest).
We now explain how varying risk aversion plays a role in thermodynamic protocols. For readers without an economics background, let us begin by recalling that in expected utility theory, the most fundamental postulate is (called the expected utility hypothesis) states that rational individuals make decisions in order to maximise their expected utility, rather than expected value [27]. In this context, risk aversion is reflected in the utility function through the concavity - which reflects the fact that a risk-averse person might prefer a guaranteed $\mathdollar 100$ over a 50 per cent chance of winning $\mathdollar 250$, even though the latter has a higher expected value. Risk-seeking tendencies are defined analogously. Hence, strategies should be chosen to reflect these risk preferences.
@@
Here, the parameter $r$ reflects the risk preferences of the gambler
  • $r>0$: Risk-aversion (concave utility) - the agent prefers guaranteed work extraction over uncertain fluctuations
  • $r=0$: Risk-neutral behaviour (linear utility, recoverable via limit $r\rightarrow 0$) - agents maximize expected work regardless of fluctuations
  • $r<0$: Risk-seeking behaviour (convex utility) - agents prefer higher amount of work extracted even if it means lower probability of success
As it is known from elementary thermodynamics [28], for a single round of work extraction in the generalised Szilard engine described above, the extracted work in terms of the outcome $x$ of the random variable $X$ can be expressed as
$w_{x}=\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)k_{B}T$ (3)
It follows using eq. 2 that the utility of outcome $x$ is
$u_{r}(w_{x})=\frac{1}{r}\bigg(1-\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (4)
By the expected utility hypothesis, a rational gambler will set their bet $Q_{X}^{A}$ as to maximise
$\sum_{x}P_{X}(x)u_{r}(w_{X}(x))=\frac{1}{r}\bigg(1-\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (5)
In appendix D, we show that for a given $r$, choice of $Q_{X}^{A}$ corresponds to
$Q_{X}^{A,r}=\frac{P_{X}(x)^{\frac{1}{1+r}}Q_{X}^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q_{X}^{B}(x^{\prime})^{\frac{r}{1+r}}}$ (6)
Where we have introduced the notation $Q_{X}^{A,r}$ to indicate that the optimiser depends on the risk aversion of the gambler. Note that as $r$ goes to zero (risk neutral) $Q_{X}^{A,r}$ goes to $P_{X}$, and as $r$ goes to $\infty$ (extreme risk aversion) $Q_{X}^{A,r}$ goes to $Q_{X}^{B}$ (not moving partition, and hence no fluctuations at all), as expected.
In the following section, we connect the strategy $Q_{X}^{A,r}$ derived above the average work extracted by a gambler with a particular level of risk aversion, as well as to Renyi divergences (the generalised free energies of the resource-theoretic approach).
III.1 Risk aversion revisited
The aim of this section is to revisit risk aversion in the thermodynamic setting and to show explicitly how the economic notion of risk aversion helps provide the missing bridge between stochastic and resource-theoretic formulations of the second law.
Recall that, in the previous section, the utility maximisation problem of the adversarial Szilard engine led to optimal strategies for Alice of the form
$Q_{X}^{A,r}=\frac{P_{X}(x)^{\frac{1}{1+r}}Q_{X}^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q_{X}^{B}(x^{\prime})^{\frac{r}{1+r}}}$ (7)
In our engine, this corresponds directly to Alice’s choice of partition placement; in economic terms, it reflects the unique optimal allocation for a CARA agent with parameter $r$. The parameter $r$ controls the risk aversion of the gambler [29] as follows:
@@ III.1 Risk aversion revisited
  • $r>0$: Risk-aversion - the agent prefers guaranteed work extraction over uncertain fluctuations. Here risk aversion increases with increasing $r$.
@@
  • $r=0$: Risk-neutral behaviour (linear utility, recoverable via limit $r\rightarrow 0$) - agents maximize expected work regardless of fluctuations.
@@
  • $r<0$: Risk-seeking behaviour - agents prefer higher amount of work extracted even if it means lower probability of success.
@@
It is useful to evaluate the expected work extracted for a particular strategy. In appendix F, we show that for risk aversion level $r$, the expected work extracted $\mathbb{E}_{r}(Q_{X}^{A})$ is given by
$\boxed{\mathbb{E}_{r}[W]=\big(\alpha D(P_{X}||Q_{X}^{B})+(1-\alpha)D_{\alpha}(P_{X}||Q_{X}^{B})\big)k_{B}T}$ (8)
where we have included the subscript $\alpha=\frac{1}{1+r}$ to connect the level of risk aversion to the Renyi parameter and to indicate that the expected value of the extracted work depends on the particular strategy and level of risk aversion.
@@
$\boxed{W_{CE}=D_{\frac{1}{1+r}}(P_{X}||Q^{B}_{X})k_{B}T}$ (9)

### v2

@@ Abstract
In thermodynamics, an agent’s ability to extract work is fundamentally constrained by their environment. Traditional frameworks struggle to capture how strategic decision-making under uncertainty — particularly an agent’s tolerance for risk — determines the trade-off between extractable work and probability of success in finite-scale experiments. Here, we develop a framework for nonequilibrium thermodynamics based on adversarial resource theories, in which work extraction is modeled as an adversarial game for an agent extracting work. Within this perspective, we consider a Szilard-type engine as a game isomorphic to Kelly gambling, an information-theoretic model of optimal betting under uncertainty — but with a thermodynamic utility function. Extending the framework to finite-size regimes, we apply a risk–reward trade-off to find an interpretation of the Rényi-divergences, in terms of extractable work for a given failure probability. By incorporating risk sensitivity via utility functions, we show that the guaranteed amount of work a rational agent would accept instead of undertaking a risky protocol is given by a Rényi divergence. This provides a unified picture of thermodynamics and gambling, and highlights how generalized free energies emerge from an adversarial setup.
@@ I Introduction
In recent decades, advances in nanotechnology and biophysics have made the study of nonequilibrium and small-scale thermodynamics increasingly important. In these regimes where the thermodynamic limit no longer applies and systems contain only finitely many interacting particles, ensemble averages fail and fluctuations dominate. Deterministic quantities such as free energy must be replaced by probabilistic, protocol-dependent notions, leading to trade-offs between extractable work and probability of success. The breakdown of determinism has motivated two main approaches to understand the second law of thermodynamics in regimes where standard assumptions break down: stochastic thermodynamics, and the resource-theoretic approach (for a review, see Ref. Sagawa (2022)).
@@
Stochastic thermodynamics Seifert (2012); Parrondo et al. (2015) embraces the inherent randomness of small, out-of-equilibrium systems by treating work as a fluctuating quantity. It describes how entropy on average must increase, while accounting for rare transient decreases. On the other hand, the resource-theoretic approach Janzing et al. (2000); Horodecki et al. (2003); Horodecki and Oppenheim (2013) reformulates the second law as a constraint on possible state transformations. It shifts the focus from averages to operational possibilities, yielding generalized second laws — expressed through Rényi divergences — that quantify how far a state is from thermal equilibrium Ruch and Mead (1976); Brandão et al. (2015). While both frameworks are profoundly successful, and their views on the second law can be related Alhambra et al. (2016), neither provides a complete, operational prescription for how an agent’s strategic choices — specifically, their tolerance for risk — directly determine the trade-off between the amount of work they can extract and the probability of successfully obtaining it in a single, finite-scale experiment. In this work, we bridge this gap by treating work extraction from the lens of expected utility theory and decision theory Von Neumann and Morgenstern (1944).
@@
There have been seminal works forging connections between stochastic thermodynamics and gambling Hirono and Hidaka (2015); Ito (2016); Manzano et al. (2020); Tohme et al. (2023); Tirone et al. (2021); Ducuara et al. (2023). One strand of research has implemented specific gambling strategies directly within thermodynamic protocols, as seen in the work extraction experiments of Manzano et al. (2020); Tohme et al. (2023). Concurrently, Ref. Ito (2016) established formal links between gambling, work extraction, and information flows from an information-thermodynamic perspective. In a separate approach, Ref. Ducuara et al. (2023) applied expected utility theory to evaluate thermodynamic processes as lotteries with predefined work payoffs. While these works successfully import concepts from one field to the other, our framework provides a deeper analogy between them and reveals a decision-theoretic aspect of thermodynamics. By formally modeling work extraction as an adversarial game implemented by a Szilard-type engine, we demonstrate that generalized free energies emerge from the necessary constraints on rational, strategic play.
@@
We develop a framework based on adversarial resource theories Arcos et al. (2025), in which an agent must choose a work-extraction protocol while contending with constraints initially imposed on the engine. This perspective casts thermodynamics as a decision-theoretic problem, where the optimal strategy is determined by the agent’s sensitivity to fluctuations. The mathematical structure of this game is formally analogous to the Kelly betting problem from information theory Kelly (1956), a connection we develop in a companion paper Arcos et al. (2025). However, a crucial distinction arises from the physical context: while financial wealth in Kelly gambling grows multiplicatively, thermodynamic work is an additive quantity. This difference dictates distinct classes of rational utility functions—Constant Relative Risk Aversion (CRRA) for gambling versus Constant Absolute Risk Aversion (CARA) for thermodynamics—which in turn shape the optimal strategies and their interpretation.
@@
The paper is structured as follows. In Section II, we formulate a Szilard-type engine as an adversarial setup between an agent (Alice) extracting the work and another agent (Bob) preparing initial constraints, and derive the average extractable work and its connection to the nonequilibrium free energy. In Section III, we analyze the risk–reward trade-off through the lens of decision theory, and in particular expected utility theory, where we identify Constant Absolute Risk Aversion (CARA) as the relevant utility class for thermodynamics. This framework allows us to compute both the certainty equivalent (the guaranteed amount of work a risk-sensitive individual would accept instead of undertaking a risky extraction protocol) as Eq. (11), and the expected value of the work extracted by a rational individual, Eq. (9) for any level of risk aversion. We find that both are parametrized by Rényi divergences. We explain why the former result has no analogue in gambling, while the latter result is related to the gambling result derived in Arcos et al. (2025).
In Section IV, we approach the problem from an information-theoretic viewpoint on the finite-size regime, where fluctuations dominate. Here, we apply results from concurrent work Arcos Enriquez (2025); Arcos et al. (2025) to show that the work extraction problem reduces to a decision-theoretic problem, introducing a fundamental risk–reward trade-off in which the work extracted if a strategy is successful can be bounded in terms of Rényi divergences $D_{\alpha}$. This gives an operational interpretation to $D_{\alpha}$ for each value of $\alpha$, in terms of the minimal work extraction given some risk tolerance. This extends the result of Dahlsten et al. (2011) who gave an interpretation of the Rényi divergences $D_{0}$ and $D_{\infty}$ in terms of bounds on extracting work in the case of an extremely risk-averse or risk-seeking agent. In both the expected utility approach and the information-theoretic approach, we give identical and explicit strategies for achieving work extraction given some risk tolerance, via Eq. (8). Together, all these results provide a unified operational interpretation that bridges stochastic and resource-theoretic views of the second law. Technical details and extended calculations are deferred to the appendices.
@@ II The adversarial Szilard engine
We begin by formalizing a thermodynamic work extraction problem as a setup between adversaries, mirroring the structure of Kelly betting, which we review in Appendix A. Consider three players: Bob, who sets initial constraints; Alice, who optimizes work extraction (analogous to a gambler allocating bets), and Charlie, a referee who enforces randomness.
We consider an empty box of volume $V$ which Bob divides into two parts by placing a partition at some position $Q^{B}$. The referee, Charlie, then samples from a binary probability distribution $P(x)$ and places a molecule on the left or right-hand side of the box according to the outcome which Alice and Bob do not know. Here, $x=0$ (11) represents that the molecule is in the left (right) side. Alice proceeds to extract work by performing isothermal compressions and expansions on the box, moving the partition to a final position $Q^{A}$ of her choice. The process is illustrated in Fig. 1. Since this setup resembles the Szilard engine Szilard (1929), we name it the adversarial Szilard engine. However, we emphasize that our setup does not involve measurement and feedback by an external “demon”, in contrast to the original Szilard’s setup.
(a) Bob places the partition at some initial position $Q^{B}$.
(b) Alice chooses a position $Q^{A}$ to which she moves the partition.
(c) When the molecule is on the left side of the box, Alice extracts work by moving it to her desired position.
(d) When the molecule is on the right side of the box, Alice must do work to move the partition to the position of her choice.
@@
Without loss of generality, assume that Bob places the partition closer to the leftmost edge of the box, so that the volume on the left-hand side can be expressed as $Q^{B}V$, and that Alice moves her partition to a position so that the volume on the left-hand side can be expressed as $Q^{A}V$ in terms of fractions $0<Q^{A}<1$ and $0<Q^{B}<1$ of the total length of the box. The work extracted by Alice depends on the molecule’s position $x=0,1$:
$w(0)=k_{B}T\ln\left(\frac{Q^{A}}{Q^{B}}\right);\ \ w(1)=k_{B}T\ln\left(\frac{1-Q^{A}}{1-Q^{B}}\right).$ (1)
@@
To maintain consistency with the notation of the literature Cover and Thomas (2006), we write $Q^{B}(0):=Q^{B}$ and $Q^{B}(1):=1-Q^{B}$, and similarly for Alice. With this notation, the amount of work extracted (which is a random variable) can be written in terms of the positions to which Alice and Bob move the partition (which are not a random variable).
The average amount of work $W$ extracted by Alice is given by
$W=k_{B}T\bigg(P(0)\ln\bigg(\frac{Q^{A}(0)}{Q^{B}(0)}\bigg)+P(1)\ln\bigg(\frac{Q^{A}(1)}{Q^{B}(1)}\bigg)\bigg),$ (2)
or equivalently,
$\boxed{W=k_{B}T\big(D(P||Q^{B})-D(P||Q^{A})\big)}$ (3)
@@
where $D(\cdot\|\cdot)$ is the standard relative entropy (the Kullback-Leibler divergence). This formula is true not only for the above binary adversarial Szilard engine (Fig. 1), but also for general multi-level engines (see Appendix B for details).
@@
The quantity (3) is maximized when Alice knows the prior $P$ and sets $Q^{A}=P$, in which case $W=k_{B}TD(P||Q^{B})$. This maximal extractable work from the system can be identified with the nonequilibrium free energy from thermodynamics up to a change in the base of the logarithm. In the particular case where Bob puts the partition in the middle, one obtains $k_{B}T(\ln 2-H(P))$, where $H(\cdot)$ is the Shannon entropy.
@@
The expression for the average extracted work by Alice is formally related to a formula appearing in the classical Kelly betting problem from information theory Kelly (1956): both involve a difference of relative entropies, and are equivalent up to an exponentiation. We develop this connection in detail in a companion paper Arcos et al. (2025). However, we note that, in Kelly’s gambling problem, wealth evolves multiplicatively, making the natural utility functions those of the CRRA type, as considered in Bleuler et al. (2020).
@@
It is also worth noting that Eq. (3) has also appeared in Ref. Kolchinsky and Wolpert (2017) in the context of work extraction with an incorrect prior distribution (see also Refs. Kolchinsky and Wolpert (2021a, b)). When Alice does not know the correct prior, our adversarial Szilard engine reduces to the formulation of Kolchinsky and Wolpert (2017) at the level of the ensemble average. We note that the ensemble average is reproduced from the empirical average in the limit of infinitely large sampling number $n\to\infty$ (assuming i.i.d.), while we will focus on the regime of finite sampling size $n$ in Section IV. Importantly, the primary focus of the subsequent arguments will be connecting Eq. (3) to expected utility theory, economics, and Kelly betting, all of which assume a known distribution in the ideal case. Therefore, we will assume that the prior distribution is known to Alice. Readers unfamiliar with Kelly betting may find it helpful to read the review in Appendix A before proceeding to the next section.
III Utility maximization: connection to expected utility formulations
@@
Recall that varying levels of risk aversion are defined in expected utility theory by the willingness to compromise between amounts of wealth and uncertainty. In this sense, the optimal strategy for different agents depends on their particular level of risk aversion. A notable example connecting the Kelly utility function to other forms of risk aversion is given in Bleuler et al. (2020), where the authors analyze Kelly gambling from the perspective of Constant Relative Risk Aversion (CRRA). CRRA describes a class of utility functions in which an individual’s relative risk aversion remains constant regardless of their level of wealth. For example, if a person has more wealth, they might risk a larger absolute amount while keeping the proportion of wealth they are willing to risk constant. This property makes CRRA particularly suitable for analyzing the Kelly paradigm, as it aligns with the proportional nature of Kelly betting, since it matches financial scenarios where wealth grows multiplicatively through compounding returns.
On the other hand, in thermodynamics, the system responds to changes in fixed energy amounts. Losing a fixed amount of energy matters the same regardless of total system energy — this is captured in the economics literature by the CARA utility function. This utility function has already proven useful in other thermodynamic contexts Ducuara et al. (2023). The distinction between the two types of risk aversion reflects how thermodynamic systems respond to additive shocks, whilst financial systems experience multiplicative shocks (where gains and losses compound like interest).
We now explain how varying risk aversion plays a role in thermodynamic protocols. For readers without an economics background, let us begin by recalling that in expected utility theory, the most fundamental postulate is (called the expected utility hypothesis) states that rational individuals make decisions in order to maximize their expected utility, rather than expected value Mas-Colell et al. (1995). In this context, risk aversion is reflected in the utility function through the concavity — which reflects the fact that a risk-averse person might prefer a guaranteed $\mathdollar 100$ over a 50 per cent chance of winning $\mathdollar 250$, even though the latter has a higher expected value. Risk-seeking tendencies are defined analogously. Hence, strategies should be chosen to reflect these risk preferences.
@@
where $w$ is the work as a random variable. The parameter $r$ reflects the risk preferences of the gambler:
@@
  • $r>0$: Risk-aversion (concave utility) — The agent prefers guaranteed work extraction over uncertain fluctuations. Here risk aversion increases with increasing $r$.
@@
  • $r=0$: Risk-neutral behavior (linear utility, recoverable via limit $r\rightarrow 0$) — The agent maximizes expected work regardless of fluctuations.
@@
  • $r<0$: Risk-seeking behavior (convex utility) — The agent prefers higher amount of work extracted even if it means lower probability of success.
For a single round of work extraction in the adversarial Szilard engine described above, the extracted work in terms of the outcome $x$ can be expressed as
$w(x)=k_{B}T\ln\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg).$ (5)
See also Appendix B for the case of multi-level engines. It follows using Eq. (4) that the utility of outcome $x$ is
$u_{r}(w(x))=\frac{1}{r}\bigg(1-\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r}\bigg).$ (6)
By the expected utility hypothesis, a rational gambler will set their bet $Q^{A}$ as to maximize
$\sum_{x}P(x)u_{r}(w(x))=\frac{1}{r}\bigg(1-\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r}\bigg).$ (7)
In Appendix C, we show that for a given $r$, the optimal choice of $Q^{A}$ corresponds to
$Q^{A,r}=\frac{P(x)^{\frac{1}{1+r}}Q^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P(x^{\prime})^{\frac{1}{1+r}}Q^{B}(x^{\prime})^{\frac{r}{1+r}}},$ (8)
where we have introduced the notation $Q^{A,r}$ to indicate that the optimizer depends on the risk aversion of the gambler. Note that as $r$ goes to zero (risk neutral) $Q^{A,r}$ goes to $P$, and as $r$ goes to $\infty$ (extreme risk aversion) $Q^{A,r}$ goes to $Q^{B}$ (not moving partition, and hence no fluctuations at all), as expected.
In our adversarial Szilard engine, the strategy $Q^{A,r}$ corresponds directly to Alice’s choice of partition placement; in economic terms, it reflects the unique optimal allocation for a CARA agent with parameter $r$. In the following, we connect $Q^{A,r}$ to the average work extracted by a gambler with a particular level of risk aversion, as well as to the Rényi divergences (the generalized free energies of the resource-theoretic approach).
@@
It is useful to evaluate the expected work extracted for a particular strategy. In Appendix E, we show that for risk aversion level $r$, the average extracted work, written as $W^{(r)}$, is given by
$\boxed{W^{(r)}=k_{B}T\left(\frac{1}{1+r}D(P||Q^{B})+\frac{r}{1+r}D_{\frac{1}{1+r}}(P||Q^{B})\right)}$ (9)
Here, we defined the Rényi divergence for $\infty<\alpha<\infty$ following to Ref. van Erven and Harremoës (2014) :
$D_{\alpha}(P||Q^{B})=\frac{1}{\alpha-1}\ln\sum_{x}P(x)^{\alpha}Q^{B}(x)^{1-\alpha}.$ (10)
Eq. (9) clarifies that the ensemble average of the extracted work depends on the particular strategy and level of risk aversion. Noticing that $\frac{1}{1+r}+\frac{r}{1+r}=1$, Eq. (9) is a weighted sum of $D(P||Q^{B})$ and $D_{\frac{1}{1+r}}(P||Q^{B})$.
@@
$\boxed{W_{CE}=k_{B}TD_{\frac{1}{1+r}}(P||Q^{B})}$ (11)
We note that the certainty equivalent for $n$ rounds is exactly $n$ times the single-round certainty equivalent due to the additivity of work and independence of rounds.